OpenStudy (anonymous):

i will meadel and fan!!!!!!!!!!!! on monday, aj read 1/6 of her book and on tuseday she read 2/6 of the book. how much did she read. A. 1/6 B. 1/5 C. 1/3 D. 1/2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

HOW DO U KNOW

OpenStudy (anonymous):

(1/6)+(2/6)=(3/6)=1/2 so D. not C as @mickyboy0202 just said

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

OHH YEAHH, I did my math wrong, sry about that. Thanks for the correct answer @tom982

OpenStudy (anonymous):

No problem, happy to help guys.